OpenStudy (nicoleced):

find (x+4) exponent of 2

OpenStudy (imtiaz7):

x^2+8x+16

OpenStudy (nicoleced):

\[(x+4)^{2}\]

OpenStudy (anonymous):

\[(x+4)^2=(x+4)(x+4)\] is a start then do the usual four multiplications

OpenStudy (nicoleced):

how? can you explain step by step? please

OpenStudy (texaschic101):

you can either use the FOIL method or you can multiply everything in the first parenthesis by everything in the second parenthesis.....so how do you want to proceed

OpenStudy (nicoleced):

but how does the foil method apply to that?

OpenStudy (texaschic101):

first of all, do you see how we got (x + 4)(x + 4) from (x + 4)^2 ?

OpenStudy (nicoleced):

oh okay i see now i was just going to ask. and then foil that, right?

OpenStudy (texaschic101):

right...do you know how ?

OpenStudy (nicoleced):

let me try it on paper

OpenStudy (texaschic101):

sure thing..

OpenStudy (nicoleced):

\[x{2} +8+4x\]

OpenStudy (imtiaz7):

here the formula is a^2+2ab+b^2 here a=x, b=4

OpenStudy (texaschic101):

no... (x + 4)(x + 4) F -- multiply first terms in each set....x * x = x^2 O -- multiply outside terms in each set....x * 4 = 4x I -- multiply inside terms in each set...4 * x = 4x L -- multiply last terms in each set...4 * 4 = 16 so now we have.. x^2 + 4x + 4x + 16 -- combine like terms x^2 + 8x + 16 any questions ?

OpenStudy (nicoleced):

\[x{2} + 16+4x + 16 ???\]

OpenStudy (nicoleced):

ahhh i see what i did wrong.

OpenStudy (nicoleced):

nope no questions... for now. thanks for your help

OpenStudy (texaschic101):

anytime :)